For any given set of probabilities, the total is always 100% or 1.0. (i.e. We know he ordered SOMETHING, so all the choices have to be accounted for.)
\(P_{total} = 1.0 = P_{tea} + P_{coffee} + P_{neither}\)
\(1.0=0.42+0.3+P_{neither}\)
\(P_{neither}= 1.0 - 0.72 = 0.28\)